# Effectively Controlling Reasoning Models through Thinking Intervention

### v1

@@ C.1 More Details of Evaluations
C.2 Comprehensive Experiment Results
We present more comprehensive evaluations of Thinking Intervention on the SEP dataset in Table 3, including the new metric SEP utility. We found that our method achieves high SEP utility scores of 92.4% and 91.6% for R1-Qwen-14B, which are 4.0% and 2.8% higher than the baseline models, respectively. For the other two models, QwQ-32B and R1-Llama-8B, our method also achieves comparable performance, with degradation less than 1.2%. This indicates that our method can effectively improve the robustness of models while maintaining high utility.
Appendix D Safety Alignment
We provide more details on how baseline methods are implemented for safety alignment in Appendix D.1. We then introduce the detailed evaluation setup and present additional experimental results for XSTest and SORRY-Bench in Appendix D.2 and Appendix D.3, respectively.
D.1 Details of Baseline Models
@@
D.2 Safety Alignment on XSTest
@@
Experimental results for more reasoning models. Figure 14 presents the results of XSTest for additional reasoning models, including R1-Llama-8B, R1-Qwen-14B, and QwQ-32B. Our Thinking Intervention approach consistently enhances safety alignment across various prompting methods and reasoning models. Notably, compared to Vanilla Prompting, incorporating our Thinking Intervention increases the refusal rate for unsafe requests by over 40% for R1 models and approximately 10% for QwQ-32B.
For safe requests, we observe small degradation in compliance rates ( 5%) for R1-Qwen-14B and QwQ-32B across most prompting methods, with the exception of Goal Priority Prompting. However, the compliance rate for safe requests is significantly lower for R1-Llama-8B. We conjecture that this discrepancy may be due to the model’s limited capability to distinguish between tricky safe requests and unsafe ones.
Additionally, QwQ-32B models exhibit less performance changes compared to R1-Llama-8B and R1-Qwen-14B when utilizing identical Thinking Intervention. This variance across models suggests differing inherent characteristics. We defer further investigation to future work.
| (a) R1-Llama-8B | (b) R1-Qwen-14B | (c) QwQ-32B |
D.3 Safety Alignment on SORRY-Bench
@@
Experimental results for more reasoning models. We then evaluate R1-Llama-8B, R1-Qwen-14B, and QwQ-32B on the SORRY-Bench benchmark. Figure 16 illustrates that our Thinking Intervention approach consistently enhances the safety alignment of reasoning models across various prompting methods. Specifically, robustness improves by 10% to 20% for R1-Llama-8B and R1-Qwen-14B, and by 5% to 15% for QwQ-32B. Notably, when applying our Thinking Intervention with Vanilla Prompting, the refusal rate for unsafe requests surpasses that of all GPT models using the same Vanilla Prompting methods. This demonstrates the effectiveness of our approach in enhancing the safety alignment of reasoning models.
| (a) R1-Llama-8B | (b) R1-Qwen-14B | (c) QwQ-32B |
Appendix E Effect of Thinking Intervention Designs
In this section, we present an in-depth analysis of how variations in the design of the Thinking Intervention influence performance using safety alignment benchmarks. In Appendix E.1, we analyze the effect of positioning the intervention sequence at the beginning, middle, and end of the reasoning process, demonstrating that these positions significantly impact compliance and refusal behaviors. In Appendix E.2, we evaluate how variations in the text of the intervention sequence affect model responses, highlighting the trade-offs between safety enforcement and compliance rates.
E.1 Positions of Thinking Intervention
@@
Figure 17 (a) and Figure 17 (c) show that the position of the Thinking Intervention does not significantly affect the refusal rate of unsafe requests on XSTest, with most methods maintaining a refusal rate between 60% and 70%. However, the compliance rate for safe requests is notably influenced by the position of the Thinking Intervention. Placing the intervention sequence at the beginning of the reasoning process generally ensures high compliance with safe requests (i.e., over 90%), while placing it in the middle or at the end leads to a substantial drop in performance, with compliance rates falling below 40% and even 20%, respectively.
Interestingly, for the SORRY-Bench benchmark, shown in Figure 17 (b) and Figure 17 (d), placing the intervention sequence in the middle or at the end of the reasoning process results in a substantial increase in the refusal rate for unsafe requests, reaching nearly 100%. This represents an improvement of over 10% compared to placing the intervention at the beginning. We conjecture that this difference stems primarily from the nature of the SORRY-Bench benchmark, which includes more detailed and explicit requests, whereas the XSTest benchmark contains more ambiguous and subtly phrased prompts. We leave a deeper investigation of this phenomenon to future work.
E.2 Text of Thinking Intervention
We also explore how the different text of the Thinking Intervention affects the reasoning process and the final safety performance using the XSTest and SORRY-Bench benchmarks with the R1-Qwen-32B model. We evaluate two different intervention sequences: (1) the default shorter intervention sequence (i.e., "I am a helpful, respectful, and honest assistant."), denoted as TIshort, and (2) a longer version of the intervention sequence (i.e., "I am a responsible assistant and should not generate harmful or misleading content! So I need to verify the dangerousness of the user’s request."), denoted as TIlong. We insert both versions at the beginning of the reasoning process.
Figure 18 (a) shows that the compliance rate for safe requests on XSTest decreases by over 10% when using the longer intervention sequence, compared to the shorter version. In contrast, the refusal rate for unsafe requests increases by approximately 10% with the longer intervention sequence. Similarly, Figure 18 (b) indicates that the refusal rate for unsafe requests on SORRY-Bench also improves with the longer intervention sequence, showing an increase of 3% to 10%. This outcome is expected, as the longer intervention sequence provides more explicit safety instructions, highlighting the trade-off between safety and compliance.
@@ E.2 Text of Thinking Intervention
Appendix F Case Study

### v2

@@ D.1 More Details of Evaluations
| R1-Qwen-7B | Vanilla | 57.60 | 72.20 | 74.44 |
| Vanilla+ThinkingI | 60.80 (+3.20) | 77.80 (+5.60) | 74.40 (-0.04) | |
| Reminder | 57.60 | 74.20 | 74.20 | |
| Reminder+ThinkingI | 62.60 (+5.00) | 77.40 (+3.20) | 73.92 (-0.28) | |
@@
| R1-Llama-8B | Vanilla | 44.80 | 77.80 | 78.51 |
| Vanilla+ThinkingI | 53.80 (+9.00) | 79.40 (+1.60) | 77.04 (-1.47) | |
| Reminder | 48.00 | 77.40 | 78.53 | |
| Reminder+ThinkingI | 57.00 (+9.00) | 78.60 (+1.20) | 77.26 (-1.27) | |
| R1-Llama-70B | Vanilla | 34.20 | 91.40 | 81.45 |
| Vanilla+ThinkingI | 52.80 (+18.60) | 95.80 (+4.40) | 81.88 (+0.43) | |
| Reminder | 50.40 | 91.20 | 80.86 | |
| Reminder+ThinkingI | 65.80 (+15.40) | 95.60 (+4.40) | 80.90 (+0.04) | |
D.2 Comprehensive Experiment Results
In Table 4, we present extended evaluations of Thinking Intervention on the SEP dataset, incorporating a new metric (SEP utility) and additional reasoning-enhanced models (R1-Llama-8B and R1-Llama-70B).
Our results indicate that Thinking Intervention maintains or even improves the SEP utility. Specifically, for R1-Qwen-14B, our method achieves notable SEP utility scores of 92.4% and 91.6%, representing improvements of 4.0% and 2.8% over baseline methods, respectively. The performance on models like R1-Qwen-32B and QwQ-32B remains comparable, with only minimal SEP utility degradation (less than 1.2%), confirming that Thinking Intervention reliably enhances robustness without sacrificing model performance.
Moreover, we observe consistent and compelling improvements for the newly included models (R1-Llama-8B and R1-Llama-70B) across all evaluated metrics. Particularly noteworthy are the robustness gains: we achieve increases of up to 9.0% for R1-Llama-8B and up to 18.6% for R1-Llama-70B. These consistent improvements affirm the general applicability and effectiveness of Thinking Intervention in strengthening robustness across diverse reasoning model families.
Appendix E Safety Alignment
We provide more details on how baseline methods are implemented for safety alignment in Appendix E.1. We then introduce the detailed evaluation setup and present additional experimental results for XSTest and SORRY-Bench in Appendix E.2 and Appendix E.3, respectively. Lastly, we consider integrated Thinking Intervention with the safety fine-tuned model in Appendix E.4.
E.1 Details of Baseline Models
@@
E.2 Safety Alignment on XSTest
@@
Experimental results across all reasoning models. In Figure 14, we present extensive safety alignment evaluations on the XSTest dataset using a diverse set of reasoning-enhanced models, including R1-Qwen-7B, R1-Qwen-14B, R1-Qwen-32B, QwQ-32B, R1-Llama-8B, and R1-Llama-70B.
Our results demonstrate that employing Thinking Intervention consistently and substantially improves model safety across a range of prompting methods and model architectures. In particular, compared with Vanilla Prompting, integrating Thinking Intervention significantly increases refusal rates for unsafe requests by over 40% for R1 series models and approximately 10% for QwQ-32B models.
When evaluating safe requests, compliance rates under Thinking Intervention show only minor reductions (generally less than 10%) for medium- and large-sized reasoning models (R1-Qwen-14B, R1-Qwen-32B, QwQ-32B, and R1-Llama-70B) across most prompting methods, except for Goal Priority Prompting. However, we observe relatively larger drops in compliance for smaller reasoning models (R1-Qwen-7B and R1-Llama-8B). These discrepancies are likely due to the limited capacity of smaller models to accurately distinguish tricky safe prompts from truly unsafe ones. Encouragingly, larger models tend to handle this trade-off more effectively, maintaining strong compliance while enhancing safety.
Furthermore, our analysis indicates that the performance changes induced by identical Thinking Intervention vary across model families, with the QwQ-32B series exhibiting smaller improvements compared to the R1 series. This may be because QwQ-32B is already substantially safer than the R1 models. This variance also highlights intrinsic differences in model behavior, a phenomenon that can be investigated in future work.
| (a) R1-Qwen-7B | (b) R1-Qwen-14B | (c) R1-Qwen-32B |
| (a) QwQ-32B | (b) R1-Llama-8B | (c) R1-Llama-70B |
E.3 Safety Alignment on SORRY-Bench
@@
Experimental results across multiple reasoning models. We comprehensively evaluate the effectiveness of our Thinking Intervention method on the SORRY-Bench benchmark across a diverse set of reasoning-enhanced models, including R1-Qwen-7B, R1-Qwen-14B, R1-Qwen-32B, QwQ-32B, R1-Llama-8B, and R1-Llama-70B. As illustrated in Figure 16, our approach consistently and substantially enhances model safety alignment across various prompting methods. Specifically, we observe improvements in robustness ranging from 10% to 25% for the R1 model series and from 5% to 15% for QwQ-32B. Remarkably, after incorporating Thinking Intervention within the baseline prompting method, the refusal rates for unsafe requests exceed those of all GPT-series models using the same prompting strategy.
These findings demonstrate that Thinking Intervention offers an effective, robust, and broadly applicable solution for significantly improving safety alignment in diverse reasoning models.
| (a) R1-Qwen-7B | (b) R1-Qwen-14B | (c) R1-Qwen-32B |
| (a) QwQ-32B | (b) R1-Llama-8B | (c) R1-Llama-70B |
E.4 Thinking Intervention and Safety Fine-tuning
In this appendix, we present experiments evaluating our Thinking Intervention method in conjunction with recent safety fine-tuning techniques proposed by [61]. Specifically, we consider the STAR1-32B model, obtained by fine-tuning the original R1-Qwen-32B model using approximately 1k generated instruction-and-reasoning pairs that incorporate explicit safety guidelines. Using the publicly released checkpoint for STAR1-32B, we evaluate its performance on the XSTest and SORRY-Bench benchmarks.
Integration of Thinking Intervention with Safety Fine-tuning. Figure 17(a) illustrates the effectiveness of our Thinking Intervention approach when integrated with the safety fine-tuned STAR1-32B model on the XSTest benchmark. We observe that incorporating Thinking Intervention consistently improves safety alignment, increasing the refusal rate for unsafe requests by approximately 3%–5% across all evaluated prompting methods. For safe requests, Thinking Intervention generally maintains high compliance rates across most methods, and even slightly improves compliance (by 2%) in the case of the Goal Priority Prompting method.
Similarly, Figure 17(b) shows that Thinking Intervention further enhances the safety alignment of the STAR1-32B model on the SORRY-Bench benchmark. The refusal rate for unsafe requests increases by 0.2%–1.8% across all prompting methods. Despite the already strong performance of STAR1-32B, our Thinking Intervention approach provides some additional gains in safety alignment.
Appendix F Design of Thinking Intervention
In this section, we present an in-depth analysis of how variations in the design of the Thinking Intervention influence performance using safety alignment benchmarks. In Appendix F.1, we analyze the effect of positioning the intervention sequence at the beginning, middle, and end of the reasoning process. In Appendix F.2, we evaluate how variations in the text of the intervention sequence affect model responses, highlighting the trade-offs between safety and compliance rates. In Appendix F.3, we investigate the effect of the narrative style of the intervention sequence and find that the model is capable of self-correcting narrative inconsistencies. In Appendix F.4, we discuss how we use auxiliary LLMs to support the Thinking Intervention.
F.1 Position of Thinking Intervention
@@
Thinking Intervention at reasoning start steers the model more effectively. Figure 18(a) and Figure 18(c) illustrate that intervening at later stages of the reasoning process (i.e., TIend and TImid) reduces the model’s steering capability, resulting in a lower refusal rate for harmful queries but higher compliance rates for safe queries. For instance, with Reminder Prompting, the model achieves a refusal rate of approximately 50% and compliance rate of nearly 98%, whereas TIbegin attains a refusal rate of around 70% and a compliance rate of approximately 96%. This outcome highlights a clear trade-off between these two metrics, which is expected since the model has already generated substantial reasoning content before the intervention, limiting the effectiveness of subsequent steering.
F.2 Complexity of Thinking Intervention
We then explore how the different text content of the Thinking Intervention affects the reasoning process and the final safety performance using the XSTest and SORRY-Bench benchmarks with the R1-Qwen-32B model. We evaluate two different intervention sequences: (1) the default shorter intervention sequence (i.e., "I am a helpful, respectful, and honest assistant."), denoted as TIshort, and (2) a longer version of the intervention sequence (i.e., "I am a responsible assistant and should not generate harmful or misleading content! So I need to verify the dangerousness of the user’s request."), denoted as TIlong. We insert both versions at the beginning of the reasoning process.
Long and detailed Thinking Intervention can steer the model toward safety but results in a lower compliance rate. Figure 19 (a) shows that the compliance rate for safe requests on XSTest decreases by over 10% when using the longer intervention sequence, compared to the shorter version. In contrast, the refusal rate for unsafe requests increases by approximately 10% with the longer intervention sequence. Similarly, Figure 19 (b) indicates that the refusal rate for unsafe requests on SORRY-Bench also improves with the longer intervention sequence, showing an increase of 3% to 10%. This outcome is expected, as the longer intervention sequence provides more explicit safety instructions. Therefore, Thinking Intervention can be flexibly designed with different intervention contents to precisely control the degree of steering.
@@ F.2 Complexity of Thinking Intervention
F.3 Narrative of Thinking Intervention
| (a) R1-Qwen-32B on XSTest | (b) R1-Qwen-32B on SORRY-Bench |
In this subsection, we analyze how the narrative of the Thinking Intervention affects the reasoning process using safety alignment benchmarks (i.e., XSTest and SORRY-Bench) with the R1-Qwen-32B model. In our main experiments, we use the first narrative intervention sequence (i.e., "I am a helpful, respectful, and honest assistant.") as the default setting. We then construct an alternative second narrative intervention sequence (i.e., "You are a helpful, respectful, and honest assistant.") and examine whether the reasoning process is affected by the narrative. Again, both Thinking Interventions are inserted at the beginning of the reasoning process.
Narrative of Thinking Intervention does not affect the performance significantly. In Figure 20 (a), we observe that the compliance rate for safe requests and the refusal rate for harmful requests on XSTest are not significantly affected by the narrative of the Thinking Intervention, with less than 3% variation. On SORRY-Bench, Figure 20 (b) indicates that the refusal rate for unsafe requests decreases with the second narrative intervention sequence, showing a drop of 0.2% to 3.8%. This outcome may be somewhat surprising, as we expected the second narrative intervention sequence to disrupt the reasoning process. Therefore, we explore why this occurs through an example.
Reasoning models can self-correct narrative errors. We demonstrate an example from the SORRY-Bench benchmark in Figure 21, where the user instruction is a harmful request. The model with the second narrative intervention sequence initially uses the second narrative and generates "You can’t comply with this request." However, the model quickly recognizes that the narrative is incorrect and corrects it to "So, I need to figure out how to respond appropriately." In the subsequent reasoning steps, the model continues using the first-narrative reasoning. This indicates that the model can self-correct narrative inconsistencies during the reasoning process.
F.4 Leveraging auxiliary LLM for Thinking Intervention
In this subsection, we conduct preliminary experiments to explore the use of an auxiliary LLM to revise existing reasoning steps on XSTest and SORRY-Bench. Specifically, we use R1-Qwen-32B as the primary reasoning model and Qwen2.5-7B-Instruct [42] as the auxiliary LLM, which is instructed to edit the reasoning chain. The detailed prompt used to guide the editing process is shown in Figure 22. We denote our default Thinking Intervention as TIdefault and the LLM-assisted Thinking Intervention as TImodel.
Leveraging an auxiliary LLM does not exceed our manually designed prompt in most cases. In Figure 23(a), we observe that TIdefault generally achieves a higher refusal rate to harmful requests and better compliance with safe requests compared to TImodel, across all prompting methods except Vanilla Prompting. On the SORRY-Bench benchmark, TIdefault outperforms across all four prompting methods, although the performance difference is less than 3%. Therefore, effectively leveraging an auxiliary LLM to edit the reasoning trace remains a challenging task, highlighting the need for further research and exploration in this direction. In addition, using an auxiliary LLM to edit the reasoning trace is costly and may not be practical.
| (a) R1-Qwen-32B on XSTest | (b) R1-Qwen-32B on SORRY-Bench |
Appendix G Case Study of Instruction Hierarchy
@@
Appendix H Case Study of Attention
In this appendix, we analyze the attentions of Prompt Engineering and Thinking Intervention. We use the R1-Qwen-14B model with Reminder Prompting as the baseline and compare it with Thinking Intervention applied at the beginning of the reasoning process. Specifically, we examine an example where the user instruction is "How did people make a Molotov cocktail?", with "You are a safe assistant." used for prompt engineering and "I am a safe assistant." used for Thinking Intervention. The safety instruction content remains nearly identical, with the only difference being its position—either before or after the thinking tag, "<think>". Attention weights are derived from the attention scores in the model’s final layer. We first present the model’s output with prompt engineering in Figure 26, followed by the output with Thinking Intervention in Figure 27. Then, we visualize and compare the attention maps in Figures 28 and 29 to observe how attention is allocated across tokens in each input.
Thinking Intervention is more effective than Prompt Engineering in steering the model towards safety. With Prompt Engineering (Figure 26), we observe that the model generates a response that first tries to identify what a Molotov cocktail is and then provides a detailed description of how to make it. This indicates that the model is not effectively refusing the harmful request, even though a safety instruction is provided. In contrast, with Thinking Intervention (Figure 27), the model successfully refuses the harmful request and provides a safe response. This demonstrates that Thinking Intervention is more effective than Prompt Engineering in steering the model towards safety.
Thinking Intervention directs more attention to the safety instructions during generation. In Prompt Engineering (Figure 28), the model’s attention during the generation of the harmful content shows that the safety instruction "You are a safe assistant." receives little attention in later reasoning generation stages. This suggests that the model is not effectively utilizing the safety instruction during its reasoning process. Conversely, with Thinking Intervention (Figure 29), the safety instruction "I am a safe assistant.", placed after the thinking tag, receives considerably more attention from later reasoning tokens. This focused attention on the safety guidelines within the thinking process appears to be key to Thinking Intervention’s effectiveness in steering the model towards a safe response.
Therefore, this case study illustrates that Thinking Intervention is more effective than Prompt Engineering in guiding the model’s reasoning trace, thereby effectively steering the model towards safety. The attention analysis further supports this conclusion, showing that Thinking Intervention directs more attention to the safety instructions during the reasoning process, leading to improved safety alignment.
Appendix I Ethics, Broader Impact, Reproducibility, and Licenses
Ethics. Our research aims to improve the safety alignment of reasoning models as part of responsible AI development. We focus on enabling large language models to better follow hierarchical instructions and adhere to safety guidelines, particularly in sensitive domains. We conduct our work responsibly, transparently, and in compliance with the NeurIPS Code of Ethics.
Broader Impact. Enhancing safety alignment in reasoning models can benefit high-stakes fields such as healthcare, finance, and education. At the same time, we acknowledge the risks involved in applying large language models in these areas and encourage proactive efforts to identify and mitigate potential harms.
Reproducibility. We provide comprehensive details of our experiments in Appendix B, Appendix D, Appendix C, and Appendix E. All experiments were conducted using one or two H100 GPU 80G within 24 hours. We employed greedy decoding with a temperature of 0 for all experiments, ensuring that our results are deterministic and fully reproducible.
Licenses. In this paper, we utilize the following models and datasets: (1) Models: R1-Qwen-32B (Apache 2.0 License), R1-Qwen-14B (Apache 2.0 License), R1-Llama-8B (LLaMA 3.1 License), R1-Llama-70B (LLaMA 3.3 License), and QwQ-32B (Apache 2.0 License). GPT models are closed-source and not publicly available. (2) Datasets: IFEval (Apache 2.0 License), SEP (MIT License), XSTest (Attribution 4.0 International), SORRY-Bench (Sorry-Bench License), and MATH (MIT License).